----------------------------------------------------------------------------- Mauna Loa Solar Observatory Observer's Log ----------------------------------------------------------------------------- Tue Nov 03 17:21:34 GMT 2026 Year:26 Doy:307 Observer: berkey Weather COMMENT BY berkey: Tue Nov 03 17:25:52 GMT 2026 MLO Current Weather Timestamp (UTC): 2026-11-03 17:23:00 Temperature : 46.17 deg F Dew Point : -1.34 deg F Pressure : 681.86 mbar Wind Speed : 5.37 mph Wind Direction : 192.9 deg Skies: clear above with patchy cirrus toward the east. We should get some data today. ___end___ GENERAL COMMENT BY berkey: Tue Nov 03 17:27:03 GMT 2026 Tower teams on site a sunrise today completing, expecting to see the erection of the rest of the tower today. ___end___ ** KCOR PROBLEM REPORTED BY berkey ** : Tue Nov 03 17:33:45 GMT 2026 Occulter edge damage seen in kcor occulter T1. These look like manufacturing issues I wonder if we are using the correct occulter (the one in the telescope is labeled T1 which can be seen in the data. ___end___ ** KCOR PROBLEM REPORTED BY berkey ** : Tue Nov 03 17:36:07 GMT 2026 Kcor started with the wrong polarization twice this morning. Third time seems to be the charm. ___end___ Tue Nov 03 17:41:09 GMT 2026 KCOR Start Synoptic Patrol BY berkey: Tue Nov 03 18:15:52 GMT 2026 We will need to close soon for dust related to some grading. ___end___ BY berkey: Tue Nov 03 18:48:33 GMT 2026 campus Construction crane is starting to move. So far light wind fro SE is keeping dusk and emissions away from the building. ___end___ Tue Nov 03 20:23:26 GMT 2026 KCOR End Patrol Tue Nov 03 20:23:27 GMT 2026 KCOR Start Calibration script: c:\kcor\mlso-calibration22deg-20171025.ini Tue Nov 03 20:38:42 GMT 2026 KCOR End Calibration Script Tue Nov 03 20:38:59 GMT 2026 KCOR Start Synoptic Patrol Tue Nov 03 20:39:00 GMT 2026 KCOR Start Synoptic Patrol GENERAL COMMENT BY berkey: Wed Nov 04 00:06:53 GMT 2026 Dome automation work done for today. Per-drilling holes for mounting equipment tomorrow. Current plan is to have normal dome access until about 10-11 and then complete the rotation upgrade. And finish the day with mounting systems for dome encoders and shutter power. ___end___ BY berkey: Wed Nov 04 00:37:18 GMT 2026 Updated the ucomp filter wheel code to made a direct HID (human interface driver) call from labview instead of pointing to a python call to interact with the HID driver. Some initial testing found that the python code was writing the wrong infomation during a filter wheel move. The wheel expected calls to be 0x80 + POS, where POS is the destination address. This is hex 80 or dec 128, but the python code was calling 80 + POS. Oddly the filter wheel mostly worked with dec 80; but this may explain some of the drop outs we had been seeing. ___end___ KCOR COMMENT BY KCOR: Wed Nov 04 01:06:21 GMT 2026 KCOR file transfer report KCor created 1205 .bin files 1205 .fts files Remote site is complete ? no rsync required. ___end___ Logger COMMENT BY berkey: Wed Nov 04 01:09:16 GMT 2026 Log closed ONSITE STAFF: ___end___ Logger COMMENT BY : Wed Nov 04 01:15:22 GMT 2026 Log closed ONSITE STAFF: ___end___